OpenStudy (anonymous):

Factor the polynomial completely: 4m^3-16m I need to check my answer!

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

ok what was your answer?

OpenStudy (anonymous):

4m(m-1)(m+4)

OpenStudy (bahrom7893):

4m^3-16m = 4m(m^2-4)=4m(m-2)(m+2)

OpenStudy (anonymous):

and how am i able to check my work @bahrom7893 ?????

OpenStudy (bahrom7893):

Because the answer I just gave you is the right one and you need to know this rule: a^2-b^2=(a-b)(a+b) sooo what is: m^2-16, which is also m^2 - 4^2